The loss due to an earthquake in a commercial building is modelled by a random
variable X with density function
f(x) = 
0.005(20 − x), for 0 < x < 20
0, elsewhere
Given that the fire loss exceeds 10, what is the probability that it exceeds 18
1
Expert's answer
2020-05-13T18:36:20-0400

We need to compute

"P(X>18|X>10)={P(X>18)\over P(X>10)}"

"P(X>10)=\displaystyle\int_{10}^{\infin}f(x)dx=\displaystyle\int_{10}^{20}0.005(20-x)dx="

"=0.005\bigg[20x-{x^2\over 2}\bigg]\begin{matrix}
 20 \\
 10
\end{matrix}=0.005(400-200-(200-50))="

"=0.25"

"P(X>18)=\displaystyle\int_{18}^{\infin}f(x)dx=\displaystyle\int_{18}^{20}0.005(20-x)dx="

"=0.005\bigg[20x-{x^2\over 2}\bigg]\begin{matrix}
 20 \\
 18
\end{matrix}=0.005(400-200-(360-162))="

"=0.01"

"P(X>18|X>10)={P(X>18)\over P(X>10)}={0.01\over 0.25}=0.04"